OpenStudy (anonymous):

Neeed help !!Use any graphing technology of your choice and answer the following questions about your observations. Please use complete sentences.

OpenStudy (anonymous):

I need help !! Use any graphing technology of your choice and answer the following questions about your observations. Please use complete sentences. Scenario 1: A. Graph y = –2x2 + 6. Describe what you see. B. –2x2 + 6 is what kind of polynomial? Classify the polynomial by degree and number of terms. C. Next, graph y = (-1/10)x2 + 6. Describe what you see. How is what you observe here different from what you observed in part A? D. If you wanted a really steep hill for your roller coaster, would you use the graph of y = –2x2 + 6 or y = (-1/10)x2 + 6? Scenario 2: A. Graph y = x3 – 2x2 + 6. Describe what you see. B. x3 – 2x2 + 6 is what kind of polynomial? Classify the polynomial by degree and number of terms. C. Next, graph y = x3 – 4x2 + 6. Describe what you see. How is what you observe here different from what you observed in part A? D. If you wanted the middle part of your roller coaster to have a steeper drop, would you use the graph of y = x3 – 2x2 + 6 or y = x3 – 4x2 + 6?